OpenStudy (anonymous):

solve the following equation . abs(3+x)+3=2

jimthompson5910 (jim_thompson5910):

abs(3+x)+3=2 abs(3+x)+3-3=2-3 abs(3+x) = -1 does that help?

OpenStudy (anonymous):

\[\Large \left| 3+x \right|+3=2\]

OpenStudy (anonymous):

means no solution ?

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

the absolute value of any number is never negative so abs(3+x) = -1 is impossible ---> no solutions for x

OpenStudy (anonymous):

thnx !

jimthompson5910 (jim_thompson5910):

np